We modified the EOM driver, so that the resonant frequency is now~ 14.75MHz. The full test will be done later.
As mentioned in PSL:1311, the resonant frequency on the EOM driver was not at 14.75MHz. Evan and I discussed about how to modify it and decided tof change L4 from 1.4uH to 3 uH, see the schematic here.
above, the driver after the inductor was replaced. The new one has a shield to reduce any magnetic field leakage. The legs are not fit with the footprint on the PCB, so I had to solder it to another wire to reach the footprint.
above: the TF of the driver measured between the drive and the mon output. Red trace shows the TF before the modification. Yellow trace shows the TF after the modification, notice the peak is at 14.75MHz, the Q is about the same.